OpenStudy (anonymous):

Students were surveyed about the number of children living in their household. The table shows the results. Write each experimental probability as a fraction in simplest form. 16. P ( one child ) ________________ 17. P ( 2 or more children ) _________________ 18. P ( at least 3 children ) ______________

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

what did you get for 16

OpenStudy (anonymous):

for what ?

jimthompson5910 (jim_thompson5910):

for #16

OpenStudy (anonymous):

oh i didnt do those questions yet i need help in them

jimthompson5910 (jim_thompson5910):

how many children are there total

jimthompson5910 (jim_thompson5910):

or how many students are there total

OpenStudy (anonymous):

33

jimthompson5910 (jim_thompson5910):

so P(one child) = (# of students who are an only child)/(# of students total) P(one child) = (11)/(33) P(one child) = 1/3

jimthompson5910 (jim_thompson5910):

do you see how to get #17 and #18 using this idea?

OpenStudy (anonymous):

ok thanks a lot :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

:)